Laniyah1234:

Volume of a Sphere = 4 3 πr3 The radius of the earth is approximately 3950 miles. The radius of Jupiter is approximately 44,400 miles. Since the radius of Jupiter is more than 11 times that of the earth, which comes closest to the volume of Jupiter ?

jhonyy9:

volume of a sphere = (4pi*r^3)/3 so any idea about this ?

jhonyy9:

note the radius of Earth ,,r" and the radius of Jupiter ,,R" given that R = 11*r volume of Earth = (4pi*r)/3 volume of Jupiter = (4pi*R)/3 rewrite this in function of ,,r" ,using R= 11*r and will get volume of Jupiter = [4pi*(11r)]/3 making them equal (4pi*r)/3 = [4pi*(11r)]/3 (4pi*r)/3 = [(4pi*r)/3)]*11 from this result that volume earth = (volume earth )*11 from this result that the volume of Jupiter is greatest by 11 time that volume of Earth hope this is understandably and helped